OpenStudy (anonymous):

how would i make an equation of a line that has a y intercept of 2 and an x intercept of 3? I need a formula I know this for a fact.

OpenStudy (anonymous):

Well you know 2 points: (0,2) and (3,0) find the slope and then pick a point to use in point slope form: y-y1=m(x-x1)

OpenStudy (anonymous):

okay so y-2=m(m-3)

OpenStudy (anonymous):

M is the slope, and you use one point in the formula. We'll use (0, 2) so x1=0 and y1=2 but we still need to find the slope so do that first

OpenStudy (anonymous):

I have no idea how to do that I havent taken algebra 2 in about 3 years so mind refreshing my memory

OpenStudy (anonymous):

i get -2/3

OpenStudy (anonymous):

Thats right. So when we plug everything in we get: y-2=-2/3(x-0) distribute the -2/3

OpenStudy (anonymous):

You should get: y-2=-2/3x add the 2 to both sides: y=-2/3x+2 and that is your answer

OpenStudy (anonymous):

yeah see here in lies the problem I only have these to choose from 1. -2X+3Y=4 2.-2X+3Y=6 3.2X+3Y=4 4.2X+3Y=6 5.3X+2Y=6

OpenStudy (anonymous):

Oh then it wants it in standard form. Move all the variables to one side

OpenStudy (anonymous):

How does that work?

OpenStudy (anonymous):

i need to get rid of that fraction first.

OpenStudy (anonymous):

Alright, sorry i havent converted to standard form in a while so i had to look it up. What you wanna do is multiply the whole equation by the lcd to get rid of the fraction. Which would be 3.

OpenStudy (anonymous):

so you get 3y=-2x+6 get the x on the same side as y 2x+3y=6

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanks now I get it

OpenStudy (anonymous):

You're welcome